# CREMA: Generalizable and Efficient Video-Language Reasoning via Multimodal Modular Fusion

### v2

@@ 3.3 Modality-Sequential and Modular Training of CREMA
Leveraging the proposed highly efficient modality-specific multi-query adapter (MMQA) and the self-gated multimodal query fusion approach, CREMA effectively processes a wide range of sensory input types without requiring tailored architecture designs. However, different input modalities possess distinct characteristics and varying quantities of information. Directly optimizing the model on diverse modality inputs may lead to suboptimal convergences, suffering from over/under-fitting on dominant/insignificant data types. To address the susceptibility of optimization when training on various modality inputs simultaneously, as shown in Figure 1 Right, we propose a simple yet effective remedy: modality-sequential modular training and adaptive early exit. Inspired by the alternating update of multimodal representation learning [75], in each iteration, we decompose the back-propagation process for different modalities in each iteration. Given the minibatch data $M=(M_{1},...,M_{n})$ containing $n$ modalities, we propagate $M$ through the model but update only the trainable modules corresponding to the target modality $M_{m}$ (i.e., the MMQA module for $m$ and the fusion module). Note that this sequential training within the minibatch helps the model remain robust to the order of modalities during training. Furthermore, we extend this to a modality-adaptive early exit strategy, allowing the model to bypass the training of a specific modality if it is considered as converged. We use the average gradient magnitude of the MMQA module weights as a metric for the early exit of each modality input. Specifically, at each training epoch $j+1$, we determine whether the model exits the training of the target modality if it satisfies the following equation: $\bar{\bm{g}}[j+1]>\tau\cdot\text{average}(\bar{\bm{g}}[:j])$ [^5] We omit the target modality index for brevity., where the temperature parameter $\tau$ and $\bar{\bm{g}}=[\bar{g}_{1},...,\bar{g}_{j+1}]$ denotes the list of the averaged gradient magnitudes obtained at the end of each training epoch. In the end, CREMA effectively learns rich information from various modalities for video-language reasoning, while mitigating unnecessary over-convergence or imbalanced training across different modalities.
@@ 4 Experiments
In this section, we first outline the overall experimental setup in Section 4.1 and show the results of the proposed CREMA on various cross-modal Video QA & reasoning tasks in Section 4.2. We further provide more insights about the training strategy, design of the fusion module, and the impact of new modalities across tasks in Section 4.3. More experiments on extra datasets (VGGSound [8] and PerceptionTest [50]), training memory comparison, the impact of sequential training, LoRA rank, MMQA initialization, and qualitative examples are included in Appendix (Section B).
@@ 4.1 Experimental Setup
Datasets & Benchmarks. We evaluate CREMA on the following video reasoning and QA tasks: SQA3D [46], MUSIC-AVQA [37], and NExT-QA [66]. We further evaluate CREMA on TouchQA and ThermalQA collected by ourselves based on public video-touch (Touch&Go [68]) and video-thermal data (Thermal-IM [60]). See Appendix (Sections A.1 and A.3) for more details.
@@
AVQA [37] V, A 80.3 60.0 77.3 74.5 77.8 76.1 81.4 68.7 69.9 64.6 67.1 70.9 73.5 18M -
LAVISH [41] V, A 85.6 65.9 81.4 80.2 81.1 80.6 84.6 69.2 78.8 65.6 69.1 73.8 76.9 21M -
BLIP-2 [38] V 86.7 58.5 76.3 87.2 93.7 90.5 81.5 72.0 81.3 64.3 70.1 74.2 78.9 108M 1.30K
@@
Pre-trained Visual Experts: We employ frozen pre-trained visual experts to extract modalities features from raw videos. Specifically, we use ZoeDepth [4], Unimatch [67], and NLL-AngMF [3] to estimate depth, flow, and normals, respectively. Modality Encoder: We use frozen modality-specific encoders to encode each modality to embedding space. We adopt ViT-G [57] for visual (frames, depth, norm, flow, touch, and thermal), BEATs [10] for audio, and follow data extraction in 3D-LLM [20] and ConceptFusion [26] for 3D point cloud. See Appendix (Section A.2) for details. Baselines & Model Implementation: We extend 3D-LLM and BLIP-2 with individual Q-Formers for each new modality as our baseline. We fully fine-tune these Q-Formers. Our Multimodal Q-Former is initialized from BLIP-2 image pre-trained one. We set $64$ LoRA rank and $32$ query tokens for all MMQA modules. See Appendix (Sections A.4 and A.5) for more details.
@@ 4.2 Main Experimental Results
MCAN [71] V 28.8 59.6 44.0 68.3 40.7 40.4 43.4 56M -
ClipBERT [34] V 30.2 60.1 38.7 63.3 42.4 42.7 43.3 135M -
ScanQA [2] P 31.6 63.8 46.2 69.5 43.8 45.3 46.5 38M -
3D-LLM [20] V 45.1 62.7 48.6 63.3 45.8 49.8 51.4 108M 1.30K
@@
LLaMA-VQA (7B) [32] V - - 72.7 - - 69.2 - - - 75.8 72.0  5M -
Mirasol3B [51] V - - - - - - - - - - 73.2 3B -
SeViLA [69] V 71.3 75.3 74.2 67.8 71.7 69.4 67.2 91.8 85.2 81.3 73.8 216M -
BLIP-2 [38] V 69.9 73.9 72.9 65.4 71.9 68.1 64.9 91.8 80.3 81.2 72.6 108M 1.30K
@@
Zero-shot Evaluation: In addition to the fine-tuning evaluation, CREMA method also achieves superior zero-shot performance on compositional video reasoning. We perform zero-shot evaluations on SQA3D and MUSIC-AVQA in Table 5. Note that Unified QA and GPT-3666We borrow the results of GPT-3 from the official technical report in [46]. with caption generated from 3D point cloud input perform well, attributed to their considerable model size and pre-trained data. We also observe that OneLLM, a universal multimodal reasoning framework equipped with Llama2-7B [62] for the LLM backbone, degenerates the performance when integrating different modalities: V, P. On the other hand, CREMA demonstrates a distinct advantage in zero-shot compositional reasoning across modalities, improving performance when combining video frames and 3D point clouds. We also compare with X-InstructBLIP (X-BLIP), a strong reasoning framework integrating various modalities with modality-specific Q-Formers as adapters to connect different encoders, and OneLLM on audio-video reasoning tasks. As shown, our method outperforms both X-BLIP (13B) and OneLLM (7.8B), by +8.1%$p$ and +10.3%$p$ (V, A), with only 4.2B parameters.
@@ 4.3 Quantitative Analysis
We note that CREMA exhibits improved compositional reasoning abilities using more diverse modalities, evident in higher average performance. The model benefits from the following crucial innovations. 1) we compel a regularization effect on the model through parameter-efficient updates, often leading to a stable and better generalization than fine-tuning large models [76, 13, 14]. In addition, the modular adaptation from the unified backbone using MMQA and the proposed novel modality fusion approach allow the model to produce compatible yet informative modality token embeddings sampled within a stable multimodal representation space. 2) Our proposed modality-sequential training and modality-adaptive early exit approach help CREMA to optimize multiple modality inputs effectively. In particular, we demonstrate the efficacy of our proposed modality-adaptive early exit strategy. Let the early stop indicator be $I_{es}=\bar{\bm{g}}[j+1]/\tau\cdot\text{average}(\bar{\bm{g}}[:j])$, implying that CREMA exits to learn the specific modality data once the corresponding $I_{es}$ reaches $1$. We observe that our adaptive early exit reduces training time/computations by $25\sim 60\%$, as demonstrated in Table 7 and Figure 2, without necessitating additional complex computations or calibration data. Furthermore, CREMA with a modality-adaptive early exit facilitates faster model convergence, which is evident in a performance increase of 1.0 $\sim$ 1.1% in both datasets, compared to CREMA trained for a similar number of epochs. Please see Section B.4 for the ablation of the modality-sequential training.
RQ3: Ablations for the modality fusion module of CREMA.
As the architectural design of the fusion module affects the fusion quality of token embeddings, we investigate our CREMA with different fusion strategies: given a concatenated multimodal token embedding $\mathbb{Q}$ obtained by a multimodal Q-Former (i.e., concat), Linear reduces the token size of $\mathbb{Q}$ through a linear projection, Mixture-of-Experts (MoE) adopts a MoE layer inside Q-Former to extract a few token embeddings, and Cross-Attention adopts extra prompts as an input and computes the cross-attention with $\mathbb{Q}$. As shown in Table 7, our proposed self-gated fusion module achieves competitive performance with Concat, while other variants decrease the average accuracy despite incorporating additional modalities besides video. Also, it performs efficiently compared to Cross-attention and Linear, requiring more computationally expensive operations to combine modality. We expect that the modality fusion via MoE or Cross-attention alters the original model architecture of the Q-former through additional parametric layers. Thus, it struggles to unlock the capabilities of those powerful designs under the parameter-efficient fine-tuning setup and the limited scale of downstream data.
@@
RQ4: The impact of new modalities on easy/hard questions.
We further delve into how adding extra modalities beyond video RGB frames (V) can enhance video reasoning problems. Following the previous work [6], which splits the dataset into the easy/hard groups based on the performance of the reference model to find subsets requiring less/more modality information, we classify question inputs in compositional video reasoning tasks (SQA3D and MUSIC-AVQA) based on the zero-shot performance of CREMA method with only V; i.e., if the model predicts correctly $\rightarrow$ easy, otherwise $\rightarrow$ hard, indicating that input examples in hard may need additional knowledge to find appropriate answers. After that, we fine-tune our CREMA and test on obtained subsets. As shown in Table 8, adding new modalities brings improvement over both easy and hard subsets. However, performance gain on the easy subset is less effective as it is already dominant to the video frame inputs, whereas information from additional modalities benefits the prediction of the hard (+4.9%p on SQA3D and +4.0%p on MUSIC-AVQA). In NExT-QA, adding new modalities marginally decreases the easy subset, but significantly boosts the hard (+8.5%p). It indicates that leveraging extra modalities can be an effective modality augmentation strategy, mitigating overfitting. Furthermore, CREMA method can be an efficient tool to determine modality importance for video reasoning benchmark designs.
@@ 5 Conclusion
Limitations & Broader Impacts. See Appendix for limitations and broader impacts discussion.
@@ Acknowledgement
We thank Jaemin Cho, Abhay Zala, Han Lin, Yi-Lin Sung, and Ziyang Wang for their valuable feedback and input for the paper. This work was supported by ARO Award W911NF2110220, DARPA KAIROS Grant FA8750-19-2-1004, ONR Grant N00014-23-1-2356, DARPA ECOLE Program No. HR00112390060, DARPA MCS Grant N66001-19-2-4031, and NSF-AI Engage Institute DRL211263. The views, opinions, and/or findings contained in this article are those of the authors and not of the funding agency.
@@
  • Additional experiments (Section B), including fine-tuning results on VGGSound&PerceptionTest (Section B.1), comparison of training memory (Section B.2), zero-shot per-task performance on MUSIC-AVQA (Section B.3), the impact of MMQA pre-training (Section B.5), the impact of LoRA rank (Section B.6), and more qualitative visualization (Section B.7).
@@ A.1 Benchmark and Dataset
We evaluate the CREMA framework on three video reasoning and QA tasks, focusing on both conventional VideoQA (requires video and language) and compositional VideoQA (requires video, language and other modalities). These include: (1) SQA3D [46]: Another compositional Video QA task, requiring the understanding of video, 3D scenes, and text. Designed for 3D situated QA, it includes 33K questions and 650 3D scenes corresponding to ego-centric videos. We apply extra depth maps to it and report results on the test part following [20]. (2) MUSIC-AVQA [37]: A compositional Video QA benchmark that involves reasoning across video, audio, and text. This dataset contains 9288 videos and 45K questions. We follow X-InsturctBLIP [1] to evaluate our CREMA and on other baselines on the high-quality real video part. We enhance it with optical flow as an extra input and report our findings on the test set. (3) NExT-QA [66]: A conventional Video QA benchmark for causal and temporal reasoning with video and text inputs. It consists of 5440 videos and 52K questions. We include optical flow, depth map, and surface normals extracted from raw videos as additional modalities. Our results are based on the validation partition following previous work [69]. (4) ThouchQA: we build a question and answering dataset based on Touch&Go [68]. As the original data only contains material class annotations, we reformulate the classification task as a QA task, by asking the model: What material is the person touching?. Our TouchQA dataset contains 714 training data and 3212 test data. (5) ThouchQA: Similar to the TouchQA dataset, we build the ThermalQA dataset on public video-thermal heatmap dataset, Thermal-IM [60]. Thermal-IM contains action labels in each video and was originally designed to predict human pose 3 seconds ago according to both video and thermal heatmap. We reformulate it as a QA task as well by asking the model: What action might have occurred before this video? and the answer is action label. Our Thermal dataset contains 1131 training data and 391 test data. (6) PerceptionTest [50]: a multimodal benchmark designed to comprehensively evaluate the perception and reasoning skills of multimodal video models. We use the multi-choice QA part of this benchmark, which contains 1955 train data and 5260 validation data. (7) VGGSound [8]: is a large-scale audio-video dataset. It contains over 200K videos with audio sounds. It was originally for audio-video classification. We formulated the classification tasks as the the open-ended QA task over 300 audio classes.

### v4

@@ 3.3 Modality-Sequential and Modular Training of CREMA
Leveraging the proposed highly efficient modality-specific multi-query adapter (MMQA) and the self-gated multimodal query fusion approach, CREMA effectively processes a wide range of sensory input types without requiring tailored architecture designs. However, different input modalities possess distinct characteristics and varying quantities of information. Directly optimizing the model on diverse modality inputs may lead to suboptimal convergences, suffering from over/under-fitting on dominant/insignificant data types. To address the susceptibility of optimization when training on various modality inputs simultaneously, as shown in Figure 1 Right, we propose a simple yet effective remedy: modality-sequential modular training and adaptive early exit. Inspired by the alternating update of multimodal representation learning (Zhang et al., 2024), we propose a sequential, modality-adaptive optimization process for each iteration. Instead of performing a joint back-propagation step for all modalities simultaneously, we decompose it into sequential modality-specific updates. This approach selectively updates the trainable weights corresponding to the target modality input, ensuring more focused and efficient learning for each modality. Specifically, given the minibatch data $M=(M_{1},...,M_{n})$ containing $n$ modalities, we propagate $M$ through the model but update only the trainable modules corresponding to the target modality $M_{m}$ (i.e., the MMQA module for $m$ and the fusion module). Note that this sequential training within the minibatch helps the model remain robust to the order of modalities during training. Furthermore, we extend this to a modality-adaptive early exit strategy, allowing the model to bypass the training of a specific modality if it is considered as converged. We use the average gradient magnitude of the MMQA module weights as a metric for the early exit of each modality input. Specifically, at each training epoch $j+1$, we determine whether the model exits the training of the target modality if it satisfies the following equation: $\bar{\bm{g}}[j+1]>\tau\cdot\text{average}(\bar{\bm{g}}[:j])$ [^5] We omit the target modality index for brevity., where the temperature parameter $\tau$ and $\bar{\bm{g}}=[\bar{g}_{1},...,\bar{g}_{j+1}]$ denotes the list of the averaged gradient magnitudes obtained at the end of each training epoch. In the end, CREMA effectively learns rich information from various modalities for video-language reasoning, while mitigating unnecessary over-convergence or imbalanced training across different modalities.
@@ 4 Experiments
In this section, we first outline the overall experimental setup in Section 4.1 and show the results of the proposed CREMA on various cross-modal Video QA & reasoning tasks in Section 4.2. We further provide more insights about the MLLM backbone, training strategy, design of the fusion module, and the impact of new modalities across tasks in Section 4.3. More experiments on extra datasets (VGGSound (Chen et al., 2020) and PerceptionTest (Pătrăucean et al., 2023)), training memory comparison, the impact of sequential training, LoRA rank, MMQA initialization, and qualitative examples are included in Appendix (Section B).
@@ 4.1 Experimental Setup
Datasets & Benchmarks. We evaluate CREMA on the following video reasoning and QA tasks: SQA3D (Ma et al., 2023), MUSIC-AVQA (Li et al., 2022), and NExT-QA (Xiao et al., 2021). We further evaluate CREMA on TouchQA and ThermalQA collected by ourselves based on public video-touch (Touch&Go (Yang et al., 2022)) and video-thermal data (Thermal-IM (Tang et al., 2023b)). See Appendix (Sections A.1 and A.3) for more details.
@@
AVQA (Li et al., 2022) V, A 80.3 60.0 77.3 74.5 77.8 76.1 81.4 68.7 69.9 64.6 67.1 70.9 73.5 18M -
LAVISH (Lin et al., 2023) V, A 85.6 65.9 81.4 80.2 81.1 80.6 84.6 69.2 78.8 65.6 69.1 73.8 76.9 21M -
BLIP-2 (Li et al., 2023b) V 86.7 58.5 76.3 87.2 93.7 90.5 81.5 72.0 81.3 64.3 70.1 74.2 78.9 108M 1.30K
@@
Pre-trained Visual Experts: We employ frozen pre-trained visual experts to extract modalities features from raw videos. Specifically, we use ZoeDepth (Bhat et al., 2023), Unimatch (Xu et al., 2023), and NLL-AngMF (Bae et al., 2021) to estimate depth, flow, and normals, respectively. Modality Encoder: We use frozen modality-specific encoders to encode each modality to embedding space. We adopt ViT-G (Sun et al., 2023b) for visual (frames, depth, norm, flow, touch, and thermal), BEATs (Chen et al., 2023a) for audio, and follow data extraction in 3D-LLM (Hong et al., 2023) and ConceptFusion (Jatavallabhula et al., 2023) for 3D point cloud. See Appendix (Section A.2) for details. Baselines & Model Implementation: We extend 3D-LLM and BLIP-2 with individual Q-Formers for each new modality as our baseline. We fully fine-tune these Q-Formers. Our Multimodal Q-Former is initialized from BLIP-2 pre-trained one. We set $64$ LoRA rank and $32$ query tokens for all MMQA modules. More details in Appendix (Sections A.4 and A.5).
@@ 4.2 Main Experimental Results
MCAN (Yu et al., 2019) V 28.8 59.6 44.0 68.3 40.7 40.4 43.4 56M -
ClipBERT (Lei et al., 2021) V 30.2 60.1 38.7 63.3 42.4 42.7 43.3 135M -
ScanQA (Azuma et al., 2022) P 31.6 63.8 46.2 69.5 43.8 45.3 46.5 38M -
3D-LLM (Hong et al., 2023) V 45.1 62.7 48.6 63.3 45.8 49.8 51.4 108M 1.30K
@@
LLaMA-VQA (7B) (Ko et al., 2023) V - - 72.7 - - 69.2 - - - 75.8 72.0  5M -
Mirasol3B (Piergiovanni et al., 2023) V - - - - - - - - - - 73.2 3B -
SeViLA (Yu et al., 2023a) V 71.3 75.3 74.2 67.8 71.7 69.4 67.2 91.8 85.2 81.3 73.8 216M -
BLIP-2 (Li et al., 2023b) V 69.9 73.9 72.9 65.4 71.9 68.1 64.9 91.8 80.3 81.2 72.6 108M 1.30K
@@
Zero-shot Evaluation: In addition to the fine-tuning evaluation, CREMA method also achieves superior zero-shot performance on compositional video reasoning. We perform zero-shot evaluations on SQA3D and MUSIC-AVQA in Table 5. Note that Unified QA (Khashabi et al., 2020) and GPT-3666We borrow the results of GPT-3 from the official technical report in (Ma et al., 2023). with caption generated from 3D point cloud input perform well, attributed to their considerable model size and pre-trained data. We also observe that OneLLM (Han et al., 2023), a universal multimodal reasoning framework equipped with Llama2-7B (Touvron et al., 2023) for the LLM backbone, degenerates the performance when integrating different modalities: V, P. On the other hand, CREMA demonstrates a distinct advantage in zero-shot compositional reasoning across modalities, improving performance when combining video frames and 3D point clouds. We also compare with X-InstructBLIP (Panagopoulou et al., 2023), a strong reasoning framework integrating various modalities with modality-specific Q-Formers as adapters to connect different encoders, and OneLLM on audio-video reasoning tasks. As shown, our method outperforms both X-BLIP (13B) and OneLLM (7.8B), by +8.1%$p$ and +10.3%$p$ (V, A).
Table 6: Comparison between CREMA with Mistra-7B and other popular Video-LLMs with comparable size LLMs. *: our reproduced.
Model (Modality) LLM   Acc.
Tr.
Params.
Video-LLaMA (V) Vicuna-7B 60.6 216M
Video-LLaVA (V) Vicuna-7B 66.3 7B
VideoChat2 (V) Vicuna-7B 68.6 $>$200M
LLaMA-VQA (V) LLaMA2-7B 72.0 5M
MotionEpic - VoT (V) Vicuna-7B 76.0 $>$100M
LLaVA-NeXT (V) Qwen1.5-7B 78.2 7B
VideoChat2* (V) Mistral-7B 78.4 $>$200M
CREMA (V, F) Mistral-7B 78.9 21M
CREMA (V, F, D) Mistral-7B 79.4 45M
CREMA with Different MLLM Backbone: We further built a new version of CREMA with VideoChat2 (Li et al., 2024) using stronger Mistral-7B as an LLM backbone and processed more frames (16 frames following the VideoChat2 setting). As shown in Table 6, we observed that our CREMA achieved impressive fine-tuning performance (79.4%) outperforming other strong video-LLMs by incorporating optical flow (F) and depth map (D) on NExT-QA while using a remarkably smaller number of trainable parameters, demonstrating the generalization ability across MLLM and efficiency of our proposed framework. We leave the CREMA with even stronger MLLMs (e.g., BLIP-3 (Xue et al., 2024), Qwen-VL (Wang et al., 2024b)) with more diverse modalities for further studies.
@@ 4.3 Quantitative Analysis
We note that CREMA exhibits improved compositional reasoning abilities using more diverse modalities, evident in higher average performance. The model benefits from the following crucial innovations. 1) we compel a regularization effect on the model through parameter-efficient updates, often leading to a stable and better generalization than fine-tuning large models (Zhao et al., 2021, Ding et al., 2022, Fu et al., 2023). In addition, the modular adaptation from the unified backbone using MMQA and the proposed novel modality fusion approach allow the model to produce compatible yet informative modality token embeddings sampled within a stable multimodal representation space. 2) Our proposed modality-sequential training and modality-adaptive early exit approach help CREMA to optimize multiple modality inputs effectively. In particular, we demonstrate the efficacy of our proposed modality-adaptive early exit strategy. Let the early stop indicator be $I_{es}=\bar{\bm{g}}[j+1]/\tau\cdot\text{average}(\bar{\bm{g}}[:j])$, implying that CREMA exits to learn the specific modality data once the corresponding $I_{es}$ reaches $1$. We observe that our adaptive early exit reduces training time/computations by $25\sim 60\%$, as demonstrated in Table 8 and Figure 2, without necessitating additional complex computations or calibration data. Furthermore, CREMA with a modality-adaptive early exit facilitates faster model convergence, which is evident in a performance increase of 1.0 $\sim$ 1.1% in both datasets, compared to CREMA trained for a similar number of epochs. Please see Section B.4 for the ablation of the modality-sequential training.
RQ3: Ablations for the modality fusion module of CREMA. As the architectural design of the fusion module affects the fusion quality of token embeddings, we investigate our CREMA with different fusion strategies: given a concatenated multimodal token embedding $\mathbb{Q}$ obtained by a multimodal Q-Former (i.e., concat), Linear reduces the token size of $\mathbb{Q}$ through a linear projection, Mixture-of-Experts (MoE) adopts a MoE layer inside Q-Former to extract a few token embeddings, and Cross-Attention adopts extra prompts as an input and computes the cross-attention with $\mathbb{Q}$. As shown in Table 8, our proposed self-gated fusion module achieves competitive performance with Concat, while other variants decrease the average accuracy despite incorporating additional modalities besides video. Also, it performs efficiently compared to Cross-attention and Linear, requiring more computationally expensive operations to combine modality. We expect that the modality fusion via MoE or Cross-attention alters the original model architecture of the Q-former through additional parametric layers. Thus, it struggles to unlock the capabilities of those powerful designs under the parameter-efficient fine-tuning setup and the limited scale of downstream data.
@@
RQ4: The impact of new modalities on easy/hard questions. We further delve into how adding extra modalities beyond video RGB frames (V) can enhance video reasoning problems. Following the previous work (Buch et al., 2022), which splits the dataset into the easy/hard groups based on the performance of the reference model to find subsets requiring less/more modality information, we classify question inputs in compositional video reasoning tasks (SQA3D and MUSIC-AVQA) based on the zero-shot performance of CREMA method with only V; i.e., if the model predicts correctly $\rightarrow$ easy, otherwise $\rightarrow$ hard, indicating that input examples in hard may need additional knowledge to find appropriate answers. After that, we fine-tune our CREMA and test on obtained subsets. As shown in Table 9, adding new modalities brings improvement over both easy and hard subsets. However, performance gain on the easy subset is less effective as it is already dominant to the video frame inputs, whereas information from additional modalities benefits the prediction of the hard (+4.9%p on SQA3D and +4.0%p on MUSIC-AVQA). In NExT-QA, adding new modalities marginally decreases the easy subset, but significantly boosts the hard (+8.5%p). It indicates that leveraging extra modalities can be an effective modality augmentation strategy, mitigating overfitting. Furthermore, CREMA method can be an efficient tool to determine modality importance for video reasoning benchmark designs.
@@ 5 Conclusion
Ethics Statement
The intended use of CREMA is to conduct video-language reasoning with the help of diverse assistance modalities. This does not have any particular potential for misuse beyond the general potential for AI technology to be used in harmful ways. Because it is based on MLLM to answer video questions, CREMA has the potential to hallucinate. Note that potential is shared widely in the MLLMs/Video-LLMs (Li et al., 2023b; 2024). The fact that CREMA answers questions with diverse modality augmentation, enhances the system performance while delimiting model overfitting thus mitigating model hallucination. This is crucial to build reliable, trustworthy video-language reasoning systems that assist educational AI.
@@ Acknowledgement
We thank the reviewers and Jaemin Cho, Abhay Zala, Han Lin, Yi-Lin Sung, and Ziyang Wang for their valuable feedback and input for the paper. This work was supported by the National Institutes of Health (NIH) under other transactions 1OT2OD038045-01, ARO Award W911NF2110220, DARPA KAIROS Grant FA8750-19-2-1004, ONR Grant N00014-23-1-2356, DARPA ECOLE Program No. HR00112390060, DARPA MCS Grant N66001-19-2-4031, and NSF-AI Engage Institute DRL211263. The views, opinions, and/or findings contained in this article are those of the authors and not of the funding agency.
@@
  • Additional experiments (Section B), including fine-tuning results on VGGSound&PerceptionTest (Section B.1), comparison of training memory (Section B.2), zero-shot per-task performance on MUSIC-AVQA (Section B.3), the impact of MMQA pre-training (Section B.5), the impact of LoRA rank (Section B.6), more qualitative visualization (Section B.7), and extra discussion on model design (Section B.8).
@@
  • License information (Section D) for the datasets, codes, and models that we used in this paper.
@@ A.1 Benchmark and Dataset
We evaluate the CREMA framework on three video reasoning and QA tasks, focusing on both conventional VideoQA (requires video and language) and compositional VideoQA (requires video, language and other modalities). These include: (1) SQA3D (Ma et al., 2023): Another compositional Video QA task, requiring the understanding of video, 3D scenes, and text. Designed for 3D situated QA, it includes 33K questions and 650 3D scenes corresponding to ego-centric videos. We apply extra depth maps to it and report results on the test part following (Hong et al., 2023). (2) MUSIC-AVQA (Li et al., 2022): A compositional Video QA benchmark that involves reasoning across video, audio, and text. This dataset contains 9288 videos and 45K questions. We follow X-InsturctBLIP (Alayrac et al., 2022) to evaluate our CREMA and on other baselines on the high-quality real video part. We enhance it with optical flow as an extra input and report our findings on the test set. (3) NExT-QA (Xiao et al., 2021): A conventional Video QA benchmark for causal and temporal reasoning with video and text inputs. It consists of 5440 videos and 52K questions. We include optical flow, depth map, and surface normals extracted from raw videos as additional modalities. Our results are based on the validation partition following previous work (Yu et al., 2023a). (4) ThouchQA: we build a question and answering dataset based on Touch&Go (Yang et al., 2022). As the original data only contains material class annotations, we reformulate the classification task as a QA task, by asking the model: What material is the person touching?. Our TouchQA dataset contains 714 training data and 3212 test data. (5) ThouchQA: Similar to the TouchQA dataset, we build the ThermalQA dataset on public video-thermal heatmap dataset, Thermal-IM (Tang et al., 2023b). Thermal-IM contains action labels in each video and was originally designed to predict human pose 3 seconds ago according to both video and thermal heatmap. We reformulate it as a QA task as well by asking the model: What action might have occurred before this video? and the answer is action label. Our Thermal dataset contains 1131 training data and 391 test data. (6) PerceptionTest (Pătrăucean et al., 2023): a multimodal benchmark designed to comprehensively evaluate the perception and reasoning skills of multimodal video models. We use the multi-choice QA part of this benchmark, which contains 1955 train data and 5260 validation data. (7) VGGSound (Chen et al., 2020): is a large-scale audio-video dataset. It contains over 200K videos with audio sounds. It was originally for audio-video classification. We formulated the classification tasks as the the open-ended QA task over 300 audio classes.